OpenStudy (anonymous):

find derivative g(x)= 6-8x^2, a=2

OpenStudy (anonymous):

Is the a-2 relevant? \[g(x) = 6-8x^2\] the derivative of say\[x^n \rightarrow nx^{n-1}\] So \(x^2 \rightarrow 2x\). does that make sense?

OpenStudy (anonymous):

oops! it's suppose to be a=2.

OpenStudy (anonymous):

Ah ok, well let's start by finding the derivative anyway. what does 8x^2 become?

OpenStudy (anonymous):

Oh yeah that was easy.

OpenStudy (anonymous):

-16 x

OpenStudy (anonymous):

so the final derivative is?

OpenStudy (anonymous):

(remember that a number is simply nx^1

OpenStudy (anonymous):

the answer is -16(2) = -32

OpenStudy (anonymous):

I had trouble with this one g(x) = 6x ^2 e^-x

OpenStudy (anonymous):

is that\[g(x) = 6x^2 e^{-x}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok use, the chain rule. Do the differential of one bit times the other + the difference of the other times the first. Does that make sense?

OpenStudy (anonymous):

Yeah, I have (12x)(e^-x)+(-e^-x)(6x^2)

OpenStudy (anonymous):

Good job

OpenStudy (anonymous):

but then I don't know how the answer turns to 6xe^-x (2-x)

OpenStudy (anonymous):

You can multiply that out and then treat the different parts as separate differentials. \[6xe^{-x}(2-x)=12xe^{-x}-6xe^{-x}\]and then differentiate them separately

OpenStudy (anonymous):

Ah I see. Thanks!!!